# Properties preserved by classes of Chu transforms

### v1

@@ Abstract.
The second author is an International Research Fellow of the Japan Society for the Promotion of Science
@@ 1. Introduction
More precisely, van Benthem introduced a class of two-sorted infinitary formulas, called the flow formulas, and proved a preservation theorem for the first-order fragment of this class under Chu transforms. In [11, Definition 2.2] Mirna Džamonja and Jouko Väänänen defined dense Chu transforms and used them to compare abstract logics by a sublogic relation. As a result, they showed that dense Chu transforms between abstract logics preserve compactness. Dense Chu transforms are relevant to other contexts as well. For example, Chu transforms between partially ordered sets automatically satisfy the density condition. This motivates the problem of characterizing which properties are preserved by dense Chu transforms.
We solve this problem in our main Theorem 4.12. This is done by isolating the class of inconsistency-flow formulas, which are flow formulas with the added capability of expressing inconsistency in the second sort. As a consequence of Theorem 4.12, we are able to characterize the properties preserved by the sublogic relation from [11, Definition 2.2], at least for their first-order fragment.
Going back to van Benthem, in [31, §7.4] he suggested the possibility of obtaining further preservation results, notably for surjective Chu transforms. We show that such a result indeed holds, as a consequence of a general interpolation theorem by Martin Otto [23, Theorem 1]. On a related note, Michael Gemignani [15] posed the problem of characterizing topological properties which are preserved by passing to a coarser topology. In Theorem 3.14 we obtain such a characterization by formulating Gemignani’s problem in the language of Chu spaces and then applying Otto’s theorem once again.
One of the main contributions of this paper is that Chu spaces and Chu transforms capture a variety of concepts in mathematics. For example, Chu transforms between partially ordered sets already appeared in the work of Jürgen Schmidt [26] in 1953, under the name of Galois correspondences of mixed type. In this paper we show that Chu transforms also encapsulate the Rudin-Keisler ordering between ultrafilters (Theorem 5.4), as well as strictly continuous functions between graphs (Proposition 6.3).
The presentation of our results is organized as follows. In Section 2 we introduce the central notions of Chu spaces and transforms, together with some of their basic properties. Section 3 contains our preservation result for surjective Chu transforms, as well as the solution to Gemignani’s problem. Section 4 is dedicated to our main Theorem 4.12 and its role in the preservation of compactness properties of abstract logics and topological spaces. Finally, our results on ultrafilters and graphs are presented in Section 5 and Section 6, respectively.
@@ 2. Chu spaces and Chu transforms: definitions and examples
Example 2.2.
(1) A typical and motivating example of a Chu space is a topological space, discussed in Subsection 2.1.
(2) Another important example are abstract logics, introduced in Subsection 2.2.
(3) In §5 we discuss partially ordered sets, represented as Chu spaces $\langle P,\leq,P\rangle$.
(4) Perhaps more surprisingly, another example is obtained in the world of ultrafilters, simply by coding an ultrafilter $U$ as a Chu space $\langle U,\supseteq,U\rangle$. This seemingly trivial representation gives us a correspondence with the Rudin-Keisler ordering, as we show in Theorem 5.4.
(5) Already a fifth example, treated in §6, consists of simple undirected graphs.
@@
The following separation properties seem to have first been considered by Barr [4, Section 6], under a slightly different terminology.
@@ A Chu space $\langle X,r,A\rangle$ is:
  • admits ${<}\kappa$-intersections if for every $B\subseteq A$ with $\lvert B\rvert<\kappa$ there exists $c\in A$ such that for all $x\in X$ we have $\langle x,c\rangle\in r\iff\forall b\in B(\langle x,b\rangle\in r)$;
  • admits ${<}\kappa$-unions if for every $B\subseteq A$ with $\lvert B\rvert<\kappa$ there exists $d\in A$ such that for all $x\in X$ we have $\langle x,d\rangle\in r\iff\exists c\in B(\langle x,c\rangle\in r)$.
@@ A Chu transform $(f,g)$ from $\langle X,r,A\rangle$ to $\langle Y,s,B\rangle$ is
In addition to [31], where they were first considered, surjective Chu transforms were defined by Marta García-Matos and Väänänen [14] in the context of abstract logics. [^3] The authors of [14] do not use the terminology of Chu spaces and transforms, but rather define the specific instance of a Chu transform between abstract logics which gives the sublogic relation, [14, Definition 2.2]. Restrictions in the second sort were used by Solomon Feferman in [13] and Otto in [23]. More recently, dense Chu transforms were introduced by Džamonja and Väänänen [11, Definition 2.2].
@@ 2.1. Topological spaces
Topological spaces can be encoded as Chu spaces. We present a slightly different angle on this by considering topological systems defined by Steven Vickers.
Definition 2.9 (Vickers [32, Definition 5.1.1]).
A topological system is a Chu space which admits ${<}\aleph_{0}$-intersections and admits ${<}\kappa$-unions for every $\kappa$.
According to Definition 2.9, if $\langle X,\tau(X)\rangle$ is a topological space then $\langle X,\in,\tau(X)\rangle$ is an extensional topological system. Let us note that $\langle X,\tau(X)\rangle$ is $T_{0}$ if and only if $\langle X,\in,\tau(X)\rangle$ is separated. For topological systems, admitting ${<}\kappa$-intersections means that the intersection of fewer than $\kappa$ many open sets is open, which true by definition if $\kappa=\aleph_{0}$ and true for larger $\kappa$ for the ${<}\kappa$-open spaces defined by Roman Sikorski in [28] and considered by David Buhagiar and Džamonja in [7].
@@
Proposition 2.10 (Vickers [32, Proposition 5.3.1]).
@@ 2.2. Abstract logics
The concept of abstract logic can be traced back to the work of Andrzej Mostowski [22], Per Lindström [18], and K. Jon Barwise [5]. For the purpose of this paper, an abstract logic is simply a Chu space $\langle M,\models,S\rangle$, where $M$ are the models, $S$ are the sentences, and $\models$ is the satisfaction relation between models and sentences.
An abstract logic $\langle M,\models,S\rangle$ is separated if and only if no two different models in $M$ are elementarily equivalent, and it is extensional if and only if no two different sentences in $S$ are logically equivalent. To admit complements means that $S$ is closed under negation and admitting ${<}\kappa$-intersections (unions) means that $S$ is closed under ${<}\kappa$-conjunctions (disjunctions).
Dense Chu transforms can be used to define the sublogic relation between abstract logics.
Definition 2.11 (Džamonja and Väänänen [11, Definition 2.2]).
Let $\langle M,\models,S\rangle$ and $\langle M^{\prime},\models^{\prime},S^{\prime}\rangle$ be abstract logics. We say that $\langle M^{\prime},\models^{\prime},S^{\prime}\rangle$ is a sublogic of $\langle M,\models,S\rangle$, in symbols $\langle M^{\prime},\models^{\prime},S^{\prime}\rangle\leq\langle M,\models,S\rangle$, if there exists a dense Chu transform from $\langle M,\models,S\rangle$ to $\langle M^{\prime},\models^{\prime},S^{\prime}\rangle$.
Further properties of abstract logics will be discussed in §4 and Subsection 5.2.
@@ 3. Flow formulas and preservation under certain types of Chu transforms
(1) The ,$\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formulas are defined recursively as follows:
  • $(v^{\mathsf{p}}_{0}=v^{\mathsf{p}}_{1})$, $(v^{\mathsf{s}}_{0}=v^{\mathsf{s}}_{1})$ and $R(v^{\mathsf{p}},v^{\mathsf{s}})$ are atomic ,$\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formulas;
  • the negation of an ,$\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formula is an ,$\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formula;
  • infinitary conjunctions and disjunctions of ,$\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formulas are ,$\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formulas;
  • if $\varphi$ is an ,$\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formula, then $\exists\bm{v}^{\mathsf{p}}\varphi$, $\forall\bm{v}^{\mathsf{p}}\varphi$, $\exists\bm{v}^{\mathsf{s}}\varphi$ and $\forall\bm{v}^{\mathsf{s}}\varphi$ are ,$\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formulas.
(2) Let $\kappa\leq\lambda$ be infinite cardinals. The fragment $\mathcal{L}^{\mathrm{II}}_{\lambda,\kappa}$ consists of formulas which contain only conjunctions and disjunctions of size $<\lambda$ and quantification over tuples of length $<\kappa$.
The interpretation of the satisfaction relation for this two-sorted language is done analogously to the classical definition for the one-sorted $\mathcal{L}_{\infty,\infty}$ language; see the book by Herbert Enderton [12, Section 4.3] for further details on many-sorted logic. In particular, we can apply usual model-theoretic techniques such as compactness and elementary extensions to the $\mathcal{L}^{\mathrm{II}}_{\omega,\omega}$ fragment.
@@
We say that an ,$\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formula $\varphi(\bm{v}^{\mathsf{p}},\bm{v}^{\mathsf{s}})$ is preserved under Chu transforms satisfying some property $P$ if, whenever $(f,g)$ is a Chu transform from $\langle X,r,A\rangle$ to $\langle Y,s,B\rangle$ and $(f,g)$ satisfies $P$, then for all $\bm{x}\in X$ and $\bm{b}\in B$ we have
@@
Note the notation $f(\bm{x})$ which denotes the tuple obtained by applying $f$ pointwise to $\bm{x}$, and similarly for $g(\bm{b})$.
@@ 3.1. Flow formulas and preservation under Chu transforms
Flow formulas form a subclass of ,$\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$ obtained by restricting equality, negation and quantification. Here are the formal definition and the first results of [31].
@@
We explain the special treatment of the equality and negation in Definition 3.3. Namely, the ,$\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-atomic formulas of the form $\lnot(v^{\mathsf{p}}_{0}=v^{\mathsf{p}}_{1})$ and $(v^{\mathsf{s}}_{0}=v^{\mathsf{s}}_{1})$ are not included in the class of flow formulas. The reason is that the main purpose of flow formulas is to characterize the properties of Chu spaces which are preserved by Chu transforms, as in Proposition 3.5. The property $\lnot(v^{\mathsf{p}}_{0}=v^{\mathsf{p}}_{1})$ is not preserved by a Chu transform $(f,g)$, unless $f$ is injective. Analogously, the property $(v^{\mathsf{s}}_{0}=v^{\mathsf{s}}_{1})$ is not preserved by $(f,g)$, unless $g$ is injective. The original setup by van Benthem did not include any equalities.
The following proposition is a generalization of [31, Proposition 3.1], where the result is established for first-order flow formulas. Since our class of formulas is larger, we provide a detailed proof of the proposition, for completeness.
@@
Using his framework without equality, van Benthem shows as in Theorem 3.6 that Proposition 3.5 is optimal, at least for the first-order fragment. Feferman [13, Theorem 1] gives another proof of Theorem 3.6, using the same framework with equality as in the present paper.
Theorem 3.6 (van Benthem [31, Theorem 6.1]).

### v3

@@ Abstract.
The author is an International Research Fellow of the Japan Society for the Promotion of Science
Mirna Džamonja was a co-author of the accepted version of this paper. She withdrew her participation as an author due to her perception of the origins of the research project and the interference in the research procedure by the originators of the project. She apologizes to the Israel Journal of Mathematics for withdrawing her name as an author of the paper, especially at the moment when she strongly feels her support to Israel
@@ 1. Introduction
More precisely, van Benthem introduced a class of two-sorted infinitary formulas, called the flow formulas, and proved a preservation theorem for the first-order fragment of this class under Chu transforms. In [13, Definition 2.2] Džamonja and Jouko Väänänen defined dense Chu transforms and used them to compare abstract logics by a sublogic relation. As a result, they showed that dense Chu transforms between abstract logics preserve compactness. Dense Chu transforms are relevant to other contexts as well: for example, Chu transforms between partially ordered sets automatically satisfy the density condition. This motivates the problem of characterizing which properties are preserved by dense Chu transforms.
We solve this problem in our main Theorem 4.14. This is done by isolating the class of inconsistency-flow formulas, which are flow formulas with the added capability of expressing inconsistency in the second sort. As a consequence of Theorem 4.14, we are able to characterize the properties preserved by the sublogic relation from [13, Definition 2.2], at least for their first-order fragment.
Going back to van Benthem, in [10, §7.4] he suggested the possibility of obtaining further preservation results, notably for surjective Chu transforms. We show that such a result indeed holds, as a consequence of a general interpolation theorem by Martin Otto [24, Theorem 1]. On a related note, Michael Gemignani [17] posed the problem of characterizing topological properties which are preserved by passing to a coarser topology. In Theorem 3.14 we obtain such a characterization by formulating Gemignani’s problem in the language of Chu spaces and then applying Otto’s theorem once again.
One of the main tenets of this paper is that Chu spaces and Chu transforms capture a variety of concepts in mathematics. For example, Chu transforms between partially ordered sets already appeared in the work of Michaël Benado [9] in 1949, under the name “connexions monotones d’espèce mixte”. In this paper we show that Chu transforms also encapsulate the Rudin-Keisler ordering between ultrafilters (Theorem 5.4) and this through an argument that generalizes a 1956 proof by Walter Rudin [32]. For more recent applications, we show that Chu transforms correspond to strictly continuous functions between graphs (Proposition 6.3), as well as adjointable maps between orthosets with 0 (Proposition 6.8).
The presentation of our results is organized as follows. In Section 2 we introduce the central notions of Chu spaces and Chu transforms and discuss some of their basic properties. Section 3 contains our preservation result for surjective Chu transforms, together with the solution to Gemignani’s problem. Section 4 is dedicated to our main Theorem 4.14 and its role in the preservation of compactness properties of abstract logics and topological spaces. Finally, our results on ultrafilters and graphs are presented in Section 5 and Section 6, respectively.
@@ 2. Chu spaces and Chu transforms: definitions and examples
Examples 2.2.
  1. (1)
    A typical and motivating example of a Chu space is a topological space, discussed in §2.1.
  2. (2)
    Another important example are abstract logics, introduced in §2.2.
  3. (3)
    In §5 we discuss partially ordered sets, represented as Chu spaces $\langle P,\leq,P\rangle$.
  4. (4)
    Perhaps more surprisingly, another example is obtained in the world of ultrafilters, simply by coding an ultrafilter $U$ as a Chu space $\langle U,\supseteq,U\rangle$. This seemingly trivial representation gives us a correspondence with the Rudin-Keisler ordering, as we show in Theorem 5.4.
  5. (5)
    Two further examples are treated in §6, namely graphs and orthosets with 0.
@@
The following separation properties seem to have first been considered by Barr [4, §6], under a slightly different terminology.
@@ A Chu space $\langle X,r,A\rangle$ is:
  • admits $\kappa-$intersections if for every $B\subseteq A$ with $\lvert B\rvert<\kappa$ there exists $c\in A$ such that for all $x\in X$ we have $\langle x,c\rangle\in r\iff\forall b\in B(\langle x,b\rangle\in r)$;
  • admits $\kappa-$unions if for every $B\subseteq A$ with $\lvert B\rvert<\kappa$ there exists $d\in A$ such that for all $x\in X$ we have $\langle x,d\rangle\in r\iff\exists c\in B(\langle x,c\rangle\in r)$.
@@
In the nineties, Jon Barwise and Jerry Seligman [8] rediscovered Chu spaces and Chu transforms as the fundamental theory that underlies information flow. More precisely, a classification $\bm{A}$, in the sense of [8, Definition 4.1], is just a Chu space $\langle\operatorname{tok}(\bm{A}),\vDash_{\bm{A}},\operatorname{typ}(\bm{A})\rangle$, where $\operatorname{tok}(\bm{A})$ is the set of tokens of $\bm{A}$, $\operatorname{typ}(\bm{A})$ is the set of types of $\bm{A}$, and $a\vDash_{\bm{A}}\alpha$ holds if and only if the token $a$ is of type $\alpha$. In addition to that, given classifications $\bm{A}$ and $\bm{B}$ together with functions $f\hat{\mkern 6.0mu}\colon\operatorname{typ}(\bm{A})\to\operatorname{typ}(\bm{B})$ and $f\check{\mkern 6.0mu}\colon\operatorname{tok}(\bm{B})\to\operatorname{tok}(\bm{A})$, we have that $(f\hat{\mkern 6.0mu},f\check{\mkern 6.0mu})$ is an infomorphism from $\bm{A}$ to $\bm{B}$, in the sense of [8, Definition 4.8], if and only if (fˇ,f^)($\check{\mkern 6.0mu},f\hat{\mkern 6.0mu}$) is a Chu transform from $\langle\operatorname{tok}(\bm{B}),\vDash_{\bm{B}},\operatorname{typ}(\bm{B})\rangle$ to $\langle\operatorname{tok}(\bm{A}),\vDash_{\bm{A}},\operatorname{typ}(\bm{A})\rangle$.
@@ A Chu transform $(f,g)$ from $\langle X,r,A\rangle$ to $\langle Y,s,B\rangle$ is
In addition to [10], where they were first considered, surjective Chu transforms were defined by Marta García-Matos and Väänänen [16] in the context of abstract logics. [^3] The authors of [16] do not use the terminology of Chu transforms, but rather define the specific instance of a Chu transform between abstract logics which gives the sublogic relation, [16, Definition 2.2]. Restrictions in the second sort were used by Solomon Feferman in [15] and Otto in [24]. Dense Chu transforms were introduced more recently by Džamonja and Väänänen [13, Definition 2.2].
@@ 2.1. Topological spaces
Topological spaces can be encoded as Chu spaces. We present a slightly different angle on this by considering the topological systems defined by Steven Vickers.
Definition 2.9 (Vickers [37, Definition 5.1.1]).
A topological system is a Chu space which admits $\aleph_{0}-$intersections and admits $\kappa-$unions for every $\kappa$.
According to Definition 2.9, if $\langle X,\tau(X)\rangle$ is a topological space then $\langle X,\in,\tau(X)\rangle$ is an extensional topological system. Let us note that $\langle X,\tau(X)\rangle$ is $T_{0}$ if and only if $\langle X,\in,\tau(X)\rangle$ is separated. For topological systems, admitting $\kappa-$intersections means that the intersection of fewer than $\kappa$ many open sets is open, which true by definition if $\kappa=\aleph_{0}$ and true for larger $\kappa$ for the $\kappa-$open spaces defined by Roman Sikorski in [35] and considered by David Buhagiar and Džamonja in [12].
@@
Proposition 2.10 (Vickers [37, Proposition 5.3.1]).
@@ 2.2. Abstract logics
[…]
An abstract logic $\langle M,\vDash,S\rangle$ is separated if and only if no two different models in $M$ are elementarily equivalent, and it is extensional if and only if no two different sentences in $S$ are logically equivalent. To admit complements means that $S$ is closed under negation and admitting $\kappa-$intersections (unions) means that $S$ is closed under $\kappa-$conjunctions (disjunctions).
A key observation is that a Chu transform $(f,g)$ from an abstract logic $\langle M,\vDash,S\rangle$ to an abstract logic $\langle M^{\prime},\vDash^{\prime},S^{\prime}\rangle$ is dense if and only if $g$ preserves satisfiability, in the sense that if $\sigma^{\prime}\in S^{\prime}$ is satisfiable then $g(\sigma^{\prime})$ must also be satisfiable. Motivated by this observation, dense Chu transforms can be used to define a sublogic relation between abstract logics.
Definition 2.11 (Džamonja and Väänänen [13, Definition 2.2]).
Let $\langle M,\vDash,S\rangle$ and $\langle M^{\prime},\vDash^{\prime},S^{\prime}\rangle$ be abstract logics. We say that $\langle M^{\prime},\vDash^{\prime},S^{\prime}\rangle$ is a sublogic of $\langle M,\vDash,S\rangle$, in symbols $\langle M^{\prime},\vDash^{\prime},S^{\prime}\rangle\leq\langle M,\vDash,S\rangle$, if there exists a dense Chu transform from $\langle M,\vDash,S\rangle$ to $\langle M^{\prime},\vDash^{\prime},S^{\prime}\rangle$.
Further properties of abstract logics will be discussed in §4.
@@ 3. Flow formulas and preservation under certain types of Chu transforms
The $\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formulas are defined recursively as follows:
  • $(v^{\mathsf{p}}_{0}=v^{\mathsf{p}}_{1})$, $(v^{\mathsf{s}}_{0}=v^{\mathsf{s}}_{1})$ and $R(v^{\mathsf{p}},v^{\mathsf{s}})$ are atomic $\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formulas;
  • the negation of an $\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formula is an $\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formula;
  • infinitary conjunctions and disjunctions of $\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formulas are $\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formulas;
  • if $\varphi$ is an $\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formula, then $\exists\bm{v}^{\mathsf{p}}\varphi$, $\forall\bm{v}^{\mathsf{p}}\varphi$, $\exists\bm{v}^{\mathsf{s}}\varphi$ and $\forall\bm{v}^{\mathsf{s}}\varphi$ are $\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formulas.
Let $\kappa\leq\lambda$ be infinite cardinals. The fragment $\mathcal{L}^{\mathrm{II}}_{\lambda,\kappa}$ consists of formulas which contain only conjunctions and disjunctions of size $\lambda$ and quantification over tuples of length $\kappa$.
The interpretation of the satisfaction relation for this two-sorted language is done analogously to the standard definition for the single-sorted $\mathcal{L}_{\infty,\infty}$ language; see the book by Herbert Enderton [14, §4.3] for further details on many-sorted logic. In particular, each two-sorted structure can be converted into a single-sorted structure with additional relation symbols to keep track of the sorts. In this way, usual model-theoretic techniques such as compactness and elementary extensions can be applied to the $\mathcal{L}^{\mathrm{II}}_{\omega,\omega}$ fragment.
@@
We say that an $\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formula $\varphi(\bm{v}^{\mathsf{p}},\bm{v}^{\mathsf{s}})$ is preserved under Chu transforms satisfying some property $P$ if, whenever $(f,g)$ is a Chu transform from $\langle X,r,A\rangle$ to $\langle Y,s,B\rangle$ and $(f,g)$ satisfies $P$, then for all $\bm{x}\in X$ and $\bm{b}\in B$ we have
@@
The notation $f(\bm{x})$ denotes the tuple obtained by applying $f$ pointwise to $\bm{x}$, and similarly for $g(\bm{b})$.
@@ 3.1. Flow formulas and preservation under Chu transforms
Flow formulas form a subclass of $\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$ obtained by restricting equality, negation and quantification. Here are the formal definition and the first results of [10].
@@
We explain the special treatment of the equality and negation in Definition 3.3. Namely, the $\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-atomic formulas of the form $\lnot(v^{\mathsf{p}}_{0}=v^{\mathsf{p}}_{1})$ and $(v^{\mathsf{s}}_{0}=v^{\mathsf{s}}_{1})$ are not included in the class of flow formulas. The reason is that the main purpose of flow formulas is to characterize the properties of Chu spaces which are preserved by Chu transforms, as in Proposition 3.5. The property $\lnot(v^{\mathsf{p}}_{0}=v^{\mathsf{p}}_{1})$ is not preserved by a Chu transform $(f,g)$, unless $f$ is injective. Analogously, the property $(v^{\mathsf{s}}_{0}=v^{\mathsf{s}}_{1})$ is not preserved by $(f,g)$, unless $g$ is injective. The original setup by van Benthem did not include any equalities.
The following proposition is a generalization of [10, Proposition 3.1], where the result is established for first-order flow formulas. Since our class of formulas is larger, we provide a detailed proof of the proposition, for completeness.
@@
Using his framework without equality, van Benthem shows as in Theorem 3.6 that Proposition 3.5 is optimal, at least for the first-order fragment. Feferman [15, Theorem 1] gives another proof of Theorem 3.6, using the same framework with equality as in the present paper.
Theorem 3.6 (van Benthem [10, Theorem 6.1]).